OpenStudy (anonymous):

In youngs double slit experiment slits are seperated by a distance d and screen is at a distance D>>d. From a slit plane if a light source of wavelength lambda (<

OpenStudy (anonymous):

@Mashy

OpenStudy (anonymous):

@experimentX

OpenStudy (anonymous):

Intensity.. argg.. i dunno what happens to intensity in double slit.. :P

OpenStudy (anonymous):

nvm. i might get it. 2 mins.

OpenStudy (anonymous):

I got 3. I dont know if its right though.

OpenStudy (anonymous):

Its good. Thank you.

OpenStudy (experimentx):

there should be a good relation between them, what is the relation between fringe width and d, D and lambda

OpenStudy (anonymous):

Well what i did is I know, I(resultant)=Imax cos^2(theta/2) where theta is path difference and according to the question I(res)=Imax/4 Now phase diff=path diff *2pi/lambda and path difference=Dy/d where y is position of the fringe. We have to find y.

OpenStudy (experimentx):

Fringe width (x) = λD/d

OpenStudy (anonymous):

I dont fringe width will be needed here.

OpenStudy (anonymous):

dont think*

OpenStudy (experimentx):

looks like I am also not getting what you are trying to say ... are you done with it. I am quite busy right now I can't focus.

OpenStudy (anonymous):

No problem. I was done with it already.

OpenStudy (experimentx):